Circular No. 5845

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA AQUILAE 1993 R. J. Ivison, University of Toronto, on behalf of the James Clerk

Maxwell Telescope nova monitoring group, reports the detection of N Aql 1993: “The flux densities on Aug. 15.3 UT were 59 ± 10, 80 ± 11, and 124 ± 15 mJy at 1.3, 1.1, and 0.8 mm, indicating a spectral index of 1.5. Earlier measurements at 1.1 mm gave 3σ < 16 mJy and 55 ± 11 mJy during July 1.4 and Aug. 8.4. The latest data suggest that cm-wave synthesis observations may soon pay dividends.”

PSR 1509–58 P. A. Caraveo, S. Mereghetti, and G. F. Bignami, Istituto di Fisica

Cosmica del CNR, Milan, communicate: “Using the New Technology Telescope at the European Southern Observatory, we have observed the field of the x-ray/γ-ray pulsar PSR 1509–58, which is also the radio pulsar with the highest observed period derivative. A likely optical counterpart (V = 22.0 ± 0.2, V–R ∼ +1.2) is clearly detected at the radio position. Accurate astrometry (rms < 0′′.5) from CCD images obtained on June 22-23 places the counterpart candidate at α = 15^{h}09^{m}59^{s}.1, δ = −58^{o}56′58′′.2 (equinox 1950.0). The object’s color is dominated by the significant reddening expected at its low galactic latitude and distance (∼ 4 kpc). If the optical emission is due to PSR 1509–58, it cannot be of thermal origin, and pulsations at the 150-ms rotation period should be detectable. In spite of the much longer spin period, the implied optical luminosity is comparable to that of the two other young pulsars, the Crab pulsar and PSR 0540–69, which have characteristic ages similar to PSR 1509–58 (1600 yr).”

SUPERNOVA 1991bl IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of an

apparent supernova (magnitude V ∼ 17.5) found on films taken 1991 July 5.93 and 10.92 UT with the OCA Schmidt telescope. SN 1991bl is located at α = 18^{h}12^{m}37^{s}.28, δ = +68^{o}58′13′′.9 (equinox 1950.0), or 2′′.7 west and 0′′.1 north of the apparent center of the galaxy (which has a brightness of m_{pg} ∼ 19 and is located at end figures 37^{s}.78, 13′′.8). Nothing is present at this location on the POSS prints or on similar films obtained on 1991 June 8 and 9 and in 1993. SN 1991bl does appear at mag ∼ 20.5 on films taken 1991 Sept. 5.88 and 4.90.

1993 August 17 Daniel W. E. Green